I need help in finding the average in 3 3/4, 2 1/4, 2, 4 1/2, 5/6, and 2/3. My problem come in when I add the bottom roll of fractions together. then what do I do?
1
Expert's answer
2012-10-16T08:02:46-0400
Let's add these numbers3 3/4, 2 1/4, 2, 4 1/2, 5/6,2/3:
3 3/4+ 2 1/4+ 2 + 4 1/2+ 5/6+2/3=11+3/4+1/4+1/2+5/6+2/3=11+(4+2)/4+(5+4)/6=11+3/2+3/2=11+3=14
To find average we should divide the sum of numbers by their amount. We have 6 numbers so average
a=14/6=7/3